OpenStudy (anonymous):

Does the rule y=4.4^x represent a linear or an exponential function? Explain

OpenStudy (anonymous):

Linear if the x-term or variable is the base. E.g. y=x+1 is a linear equation Exponential if the x-term or variable is in the exponent. E.g. y=4^x or y=3.4^2x

OpenStudy (anonymous):

exponential for a linear function is of the form y=ax+b